A group of researchers wants to estimate the true mean skidding distance along a new road in a certain forest. The skidding distances (in meters) were measured at 20 randomly selected road sites. These values are given in the accompanying table. Complete parts a through d.

487

350

458

196

289

407

571

438

549

382

300

435

184

260

275

405

313

311

141

426

a. Estimate the true mean skidding distance for the road with a 95 % confidence interval.,

(___,___) (Round to one decimal place as needed.)

b. Give a practical interpretation of the interval, part a.

The confidence interval means that we are ______% confident that the true

a ) mean skidding distance

B) skidding distance standard deviation

C) skidding distance variance

for

A) the sampled points

B) the whole road

C) is in the interval.

c. What conditions are required for the inference, part b, to be valid? Are these conditions reasonably satisfied? Check all correct statements below.

A. The population distribution must be approximately normal.

B The sample size must be much smaller than the population size.

C The sample must be selected carefully to represent the whole population.

D. The sample must be randomly selected from the population.

Is it reasonable to assume that the population distribution is approximately normal?

A. Yes, because the sample distribution is approximately normal.

B. No, because the sample distribution is definitely not normal.

C. The condition does not apply.

Is it reasonable to assume that the sample is random?

A. Yes, because the problem statement indicates that it is.

B. The condition does not apply.

C. No, because we do not know the methods used by the researchers.

Is it reasonable to assume that the sample size is much smaller than the population size?

A.Yes, because there are many points that could be chosen along the road.

B. No, because the length of the road is unknown.

C. The condition does not apply.

Is it reasonable to assume that the sample was selected carefully to represent the whole population?

A. Yes, because the problem statement indicates that it was.

B. No, because we do not know the methods used by the researchers.

C. The condition does not apply.

d. A logger working on the road claims the mean skidding distance is at least 432

meters. Do you agree?

a) No

b)Yes

, because of the value 432 lies

a) inside

b )outside

the confidence interval.
Answer #1

(a)

From the given data,the following statistics are calculated:

n = 20

ar{x} =7177/20 = 358.85

s = 117.6839

SE = s/sqrt{n}

= 117.6839/V20 = 26.3149

alpha = 0.05

ndf = 20 - 1 = 19

From Table, critical values of t = pm 2.0930

So,

Confidence interval:

358.85 pm (2.0930 X 26.3149)

= 358.85 pm 55.0771

= (303.8, 413.9)

(b)

The confidence level means that we are 95% confident that that the true (a) mean skidding distance (C) is in the interval.

(c)

Correct options:

A. The population distribution should be approximately normal.

B. The sample size should be smaller than the population size.

C. The sample must be selected carefully to represent the whole population.

D. The sample must be randomly selected from the population.

(i) Correct option:

C.   The condition does not apply.

(ii) Correct option:

A. Yes, because the problem statement indicates that it is.

(iii) Correct option:

A. Yes, there are many points that could be chosen along the road.

(iv) Correct option:

A. Yes, because the problem statement indicates that it was.

(v) Correct option:

(a) No

, because the value 432 lies (b) outside the confidence interval.